### v1

# Temporal and Semantic Evaluation Metrics for Foundation Models in Post-Hoc Analysis of Robotic Sub-tasks

Abstract
Recent works in Task and Motion Planning (TAMP) show that training control policies on language-supervised robot trajectories with quality labeled data markedly improves agent task success rates. However, the scarcity of such data presents a significant hurdle to extending these methods to general use cases. To address this concern, we present an automated framework to decompose trajectory data into temporally bounded and natural language-based descriptive sub-tasks by leveraging recent prompting strategies for Foundation Models (FMs) including both Large Language Models (LLMs) and Vision Language Models (VLMs). Our framework provides both time-based and language-based descriptions for lower-level sub-tasks that comprise full trajectories. To rigorously evaluate the quality of our automatic labeling framework, we contribute an algorithm SIMILARITY to produce two novel metrics, temporal similarity and semantic similarity. The metrics measure the temporal alignment and semantic fidelity of language descriptions between two sub-task decompositions, namely an FM sub-task decomposition prediction and a ground-truth sub-task decomposition. We present scores for temporal similarity and semantic similarity above 90%, compared to 30% of a randomized baseline, for multiple robotic environments, demonstrating the effectiveness of our proposed framework. Our results enable building diverse, large-scale, language-supervised datasets for improved robotic TAMP.
I Introduction
Training robotic agents on language-supervised robot trajectory data generally improves their success rates for Task and Motion Planning (TAMP) applications [1, 2]. However, the scarcity of large-scale, language-annotated robot trajectory data, particularly at the granular sub-task level, hinders mass adoption of these methods. The lack of clean, diverse, large-scale and accessible datasets also hampers the creation of Foundation Models (FMs) [3] tailored to robotic use cases [2, 4]. This prevents robotic researchers from reaching an “ImageNet Moment” [5]: a moment which revolutionized the field of computer vision as a result of creating a diverse and vast dataset of labeled images.
Recent works tackle the issue of data scarcity by focusing their efforts on basic hurdles: manually collecting, cleaning, and curating new or existing datasets into massive, open-source repositories [1, 2, 4]. However, efforts to scale-up language annotations, despite their well-known benefits, are often neglected. Prior works that offer automated language annotation methods predominantly generate high-level language instructions for full trajectories [6] or rely on computationally expensive validation methods [7].
Figure 1: Our approach evaluates a Foundation Model’s (FM) ability to temporally and semantically decompose a robot trajectory into a sub-task decomposition. We compare an FM sub-task decomposition, $\hat{\mathcal{S}}$, with a ground-truth, human labeled sub-task decomposition, $\mathcal{S}$, through our core contributions of temporal and semantic alignment metrics.
This paper proposes an alternative framework, shown in Fig. 1, which leverages powerful FMs including Visual Language Models (VLMs) and Large Language Models (LLMs) to automatically generate natural language descriptions and annotations for robot trajectories at sub-task level. In this paper, the VLMs we use include GPT-4V [8], Gemini Pro Vision, [9], and Video-LLaVa [10], and the LLM we use is Gemini Pro. Compared to traditional manual annotation via platforms like MTurk [11], our approach offers significant advantages: scalability and cost-effectiveness. More concretely, a human annotator earning the median wage, \$14.41 per hour [12], would need to annotate full trajectories in less than 30 seconds in order to be as cost effective as our method with GPT-4V. Additionally, our method allows for incorporating rich, nuanced language at a granular sub-task level into robot instructions, improving the capabilities of robotics and Embodied AI agents [13] in understanding a wider range of instructions for Task and Motion Planning. Lastly, our framework is both hardware and simulation environment agnostic.
To ensure the quality of FM-generated annotations, we define two metrics: temporal similarity, $\tau_{k}$, and semantic similarity, $\tau_{\zeta}$. Both metrics compare ground-truth annotations to FM-generated annotations, quantifying their alignment and effectiveness. These metrics allow researchers to assess the quality of the proposed framework when applied to novel tasks. This methodology only requires a test set, i.e, a small sample of labeled data, to evaluate the general performance of the FM-driven framework. Collecting a test set is far less burdensome than collecting a dataset for finetuning. Our experimental results, accumulated over 200 unique robot trajectories with 2400 variations of FM calls, demonstrate the proposed framework produces valid annotations with a $\tau_{k}$ and $\tau_{\zeta}$ both above 90% for contextually informative prompts compared to a randomized baseline near 30%, indicating our method’s effectiveness in adhering to the semantic and temporal realities of the ground-truth.
Our core contributions are:
  • A framework to automate trajectory annotations using an FM to ingest raw, multi-modal trajectory data and predict a sub-task decomposition, $\hat{\mathcal{S}}$.
  • Similarity metrics, $\tau_{k}$ and $\tau_{\zeta}$, to quantitatively evaluate the temporal and semantic alignment between two sub-task decompositions: an actual $\mathcal{S}$ and prediction $\hat{\mathcal{S}}$.
  • Results of the metrics applied across concrete examples, showcasing the efficacy of FMs in decomposing complex, multi-modal data.
II Related Work
Creating Natural Language Annotations for Robot Trajectories: Prior works generally rely on human annotators to create robot trajectories and their corresponding language descriptions [14, 15, 16, 17, 1, 18, 2]. Our focus is on developing automated methods to generate this data.
There are two works which closely relate to our proposed framework. First, Ha et al. [7] develop a framework that iteratively queries LLMs to break down high-level task instructions into sub-tasks with natural language (NL) descriptions. These subtasks are then grounded in 6-DoF action primitives which create executable trajectories. The trajectory data is vetted via simulation which provides either a success or failure signal. The failure signal initiates a re-querying of the LLM to restart the search for valid trajectory data. We find our method complementary to Ha et al.’s [7] framework, as our method relies on ground-truths generated by a Finite State Machine (FSM) controller, reducing the need for expensive re-queries to the FM. Additionally, our novel similarity metrics can be introduced prior to Ha et al.’s [7] primitive grounding step, further reducing the need for FM re-queries. The scores provided by these metrics can indicate FM response quality without running the full trajectory through a simulation. Moreover, semantic alignment, $\tau_{\zeta}$, provides finer-grain measurement over language quality than task performance metrics, which can miss nuance and noise in the synthetically generated language data. Second, Xiao et al. [6] leverage VLMs to provide automatic NL annotations to robot trajectories; but their methodology can be expensive, requiring roughly 5,600 human-labeled trajectories to fine-tune an off-the-shelf model to produce annotations. By contrast, we use prompting and rely on state-of-the-art FMs’ reasoning and in-context-learning capabilities to produce the annotations. Our approach only leverages human annotations for evaluating the quality of FM-produced annotations via our proposed SIMILARITY algorithm. Lastly, our approach produces annotations at sub-task level whereas Xiao et al. [6] provide high-level descriptions for entire trajectories.
Alternative methods at the intersection of task planning and Explainable AI (XAI) [19] leverage FMs to provide annotations for TAMP. These works are often too fine-grained, relying on the granularity of ROS 2 log files [20]. These works are also often too high-level, providing high-level summaries of trajectories [21] that cannot be used as training data. While explainability is not the primary goal of our work, we recognize this as a future application of our method.
Decomposing Trajectories: Raj et al. [22] treat trajectory decomposition as semantic parsing and are similarly focused on detecting sub-tasks to train a detection algorithm with labeled video clips and language annotation pairs. Prior works in computer vision focus on similar concepts to alleviate manual annotation of video segments, [23], [24], [25]. These methods rely on training detection algorithms rather than prompting multi-modal FMs.
FMs + Task and Motion Planning: Liu et al. [26] combines the strengths of classical planners with modern LLMs to develop long-horizon optimal plans from PDDL queries. Singh et al. [27] prompt LLMs with program-like specifications of the available actions primitives, objects in an environment, and executable, example programs to produce plans. Ahn et al. [28] ground the LLM-generated plans by multiplying each candidate action’s probability under FLAN [29] with the action’s value function, which serves as a proxy for affordance. Huang et al. [30] leverage Chain-of-Thought prompting techniques [31] to produce LLM-based plans. While there is an ever growing literature of FM-based planners [32, 33, 34], our method provides data which supports these applications. Put simply, we perform post-processing of trajectories to develop language-supervised sub-tasks to be used as training data. We do not directly target FM-based planning for TAMP.
III Problem Statement
III-A Preliminaries and Definitions
A trajectory, $T$, is a complete sequence of robot and non-robot states, $x\in X\subset\mathbb{R}^{p}$, control inputs, $u:X\mapsto U\subset\mathbb{R}^{q}$, and 3 channel images, $\mu\in\mathcal{I}\subset\mathbb{R}^{H\times W\times 3}$, that describes the evolution of robot states and non-robot states over all time steps. Robot states may encompass the x-y-z positions of a robot’s end-effector, whereas non-robot states may include the x-y-z positions of external objects. We assume a trajectory is propagated forward by discrete-time dynamics with step, $k\in[0,K]\subset\mathbb{N}^{+}$, as:
$x_{k+1}=f(x_{k},u_{k}),$ (1)
where $f:\mathbb{R}^{p}\times\mathbb{R}^{q}\mapsto\mathbb{R}^{p}$ serves as the dynamics for a robotic agent within a particular environment. The $k$ in $x_{k}$, $u_{k}$, and $\mu_{k}$ denotes the signal at step $k$. We denote the trajectory “length” as the total number of steps, $K$. While $\mu_{k}$ is not directly in Eq. 1, $\mu_{k}$ captures a visual representation of the environment. We consider the control policy to be a high-level TAMP controller, for example a Finite State Machine (FSM) or a Behavior Tree (BT) [35], [36].
Definition 1 (Trajectory Data)
A tuple of state, control, and an image recorded at every step, $k$, that is generated by a trajectory is defined as:
$\mathcal{D}=\{(x_{k},u_{k},\mu_{k})\}_{k=0}^{K}.$
We denote $\mathcal{D}^{kxu}\subset\mathcal{D}$ as textual data, which only contains $\{(x_{k},u_{k})\}_{k=0}^{K}$, and denote $\mathcal{D}^{\mu}\subset\mathcal{D}$ as visual data, which only contains $\{\mu_{k}\}_{k=0}^{K}$, so that $\mathcal{D}^{kxu}\cup\mathcal{D}^{\mu}=\mathcal{D}$.
The FM’s prompt, P, is the entire input to the FM, which incorporates $\mathcal{D}$ and context, C. We will use the description for P and C explained in [37]. The C section that provides the task to the FM is denoted herein as $^{task}\subset\textbf{C}$ C. The C section that describes the data is denoted as $^{\mathcal{D}}\subset\textbf{C}$ C. Importantly, a section of C can provide a One-Shot example, denoted herein as $^{1S}\subset\textbf{C}$ C. Zero examples, Zero-Shot, is denoted herein as $^{0S}\subset\textbf{C}$ C, and of course, $\textbf{C}^{1S}$ and $\textbf{C}^{0S}$ are mutually exclusive in C.
Definition 2 (Sub-task)
An atomic and complete motion or process, bounded temporally and concisely described by a sentence, that collectively contributes to accomplishing a larger, complex goal is defined as:
$s=(k_{\text{start}},k_{\text{end}},\zeta).$
A sub-task, $s$, includes a start step, $k_{\text{start}}$, end step, $k_{\text{end}}$, and a natural language description, $\zeta$. Both $k_{\text{start}}$ and $k_{\text{end}}$ are the same step type as in Definition 1. They delineate the temporal bounds of the sub-task, establishing its commencement and completion within $T$. The sub-task description, $\zeta$, provides a concise, semantic natural language description.
Definition 3 (Sub-task Decomposition)
A temporally ordered collection of sub-tasks, where each sub-task is indexed by $n\in[0,N]\subset\mathbb{N}^{+}$, defined as:
$\mathcal{S}=\{(k_{\text{start}},k_{\text{end}},\zeta)\}_{n=0}^{N}=\{s_{n}\}_{n=0}^{N}.$
We denote $K_{\mathcal{S}}$ and $N_{\mathcal{S}}$ as the number of steps and number of sub-tasks, respectively, in $\mathcal{S}$. Note that the map from $T$ to $\mathcal{S}$ is not injective. Practically, this means an $\mathcal{S}$ for a given $T$ is determined by a designer who chooses language descriptors and the temporal partitions’ granularity.
III-B Robotics Manipulator Example
Consider a robotic environment Stack that consists of a manipulator picking up Cube A, stacking Cube A on Cube B, then returning to a home position. Each state at a step, $x_{k}$, consists of the robot’s x-y-z end-effector position and the non-robot states, which here are each cubes’ x-y-z position and the end-effector x-y-z relative position to each cubes’ x-y-z position. The control signal at each step, $u_{k}$, is the desired manipulator joint pose. The image at each step, $\mu_{k}$, is a $256\times 256$ RGB image. The task planner is a Finite State Machine, which carefully sequences sub-tasks together to achieve the goal of stacking Cube A on Cube B.
Let us assume the entire trajectory length is 62, so that $K_{\mathcal{S}}=62$, and the number of sub-tasks is 8, so that $N_{\mathcal{S}}=8$. After $\mathcal{D}$ is recorded, an entire sub-task decomposition, $\mathcal{S}$, can be written as:
[…]
$\dots$
[…]
We will denote this $\mathcal{S}$ as the “ground-truth” because the trajectory evolved from the FSM and therefore the resulting sub-task decomposition is imposed by the FSM; however, there can be multiple ground-truth $\mathcal{S}$ for a given trajectory. This entire $\mathcal{S}$ is shown in the left column in Table I.
III-C Main Problems
Problem 1 (Steer an FM to Label a Trajectory)
Design a prompt, P, incorporating both C and $\mathcal{D}$, for an FM to output a prediction sub-task decomposition, $\hat{\mathcal{S}}$, that closely matches a ground-truth sub-task decomposition.
The FM is used “as is” with no retraining or fine-tuning to accurately form a sub-task decomposition prediction, $\hat{\mathcal{S}}$, that aligns closely with ground-truth, $\mathcal{S}$. This method capitalizes on the FM’s pre-trained knowledge base, allowing it to apply existing patterns and concepts to new, unseen tasks. This approach enhances the efficiency and scalability of sub-task decomposition through the FM’s applicability for automating nuanced, task-specific process definitions.
Problem 2 (Evaluate the FM Prediction)
Quantitatively assess the prediction quality, where the core of the evaluation focuses on temporal and semantic alignment between ground-truth $\mathcal{S}$ and FM prediction $\hat{\mathcal{S}}$.
Temporal alignment ensures $\hat{\mathcal{S}}$ is the correct sequence and duration reflecting the dynamics of the task accurately. Meanwhile, semantic alignment evaluates the relevance and appropriateness of the language descriptions assigned to each sub-task, ensuring they accurately describe the underlying action or intention. Together, these measures comprehensively gauge the FM prediction effectiveness, ensuring they are accurate in timing and contextually meaningful.
IV Approach
IV-A FMs for Auto-Decomposition of Robotic Sub-tasks
A main challenge lies in crafting a prompt, P, with context, C, that unlocks the balance between an FM’s “parametric” knowledge [38], analysis capability, and translation from P to $\hat{\mathcal{S}}$. Below is our method for building such a P.
The FM’s task and response template is assigned through $\textbf{C}^{task}$, shown as TASK_DESCRIPTION in Fig. 2. The trajectory data description, $\textbf{C}^{\mathcal{D}}$, briefly describes the expected $\mathcal{D}$ schema, shown as TEXT_DATA_DESCRIPTION and VIDEO_DATA_DESCRIPTION in Fig. 2. The $\mathcal{D}$ is placed in its entirety in the P.
The core C piece shown in Section V to enable high prediction quality is the “One-Shot” example, $\textbf{C}^{1S}$, shown by IN_CONTEXT_EXAMPLE in Fig. 2. The $\textbf{C}^{1S}$ is actually a snippet of $\mathcal{D}^{kxu}$ (from a hold-out $\mathcal{D}$, $\mathcal{S}$ pair) for each environment to provide only a partial example. When IN_CONTEXT_EXAMPLE is omitted, we consider the C to include $\textbf{C}^{0S}$ rather than $\textbf{C}^{1S}$.
To summarize, $\textbf{C}^{task}$ sets the FM’s task and provides a response template, $\textbf{C}^{\mathcal{D}}$ briefly describes the $\mathcal{D}$, $\textbf{C}^{1S}$ provides a partial example, and $\mathcal{D}$ provides two input modalities through $\mathcal{D}^{kxu}$ and $\mathcal{D}^{\mu}$.
Figure 2: A Prompt, P, with context, C, and Trajectory Data, $\mathcal{D}$, as described in Section IV-A. The key sections, in blue bold text, represent the different aspects of the P, including the FM tasking, $\textbf{C}^{task}$; partial in-context example, $\textbf{C}^{1S}$; textual data, $\mathcal{D}^{kxu}$; and visual data, $\mathcal{D}^{\mu}$.
IV-B Similarity Metrics Formulation
We propose SIMILARITY in Algorithm 1, an algorithm that calculates metrics to address 2 and compare the difference between $\mathcal{S}$ and $\hat{\mathcal{S}}$ through a temporal score, $\tau_{k}$, and a semantic score, $\tau_{\zeta}$. SIMILARITY attributes include:
  • $\tau_{k}\in[0,1]$ and $\tau_{\zeta}\in[-1,1]$, where 1 in both scoring metrics occurs only when sub-task decompositions are equal.
  • Symmetry, such that the order of the arguments does not affect the result, i.e., $\texttt{SIMILARITY}(\mathcal{S},\hat{\mathcal{S}})=\texttt{SIMILARITY}(\hat{\mathcal{S}},\mathcal{S})$.
  • Flexibility in semantic similarity by utilizing any pre-trained text encoding function.
  • The ability to handle two inputs, $\mathcal{S}$ and $\hat{\mathcal{S}}$, that may contain a different number of sub-tasks, $N_{\mathcal{S}}\neq N_{\hat{\mathcal{S}}}$, and different number of steps, $K_{\mathcal{S}}\neq K_{\hat{\mathcal{S}}}$.
Brief Algorithm Description: The algorithm compares each $s_{n}$ in $\mathcal{S}$ to each $\hat{s}_{n}$ in $\hat{\mathcal{S}}$. Sub-tasks temporally intersect when the start and end indices of $s_{n}$ are within the start and end indices of any $\hat{s}_{n}$, implying an Intersection over Union (IOU) is greater than zero. For all temporally intersecting sub-tasks the algorithm determines a temporal score, $\tau_{k}$, based on IOU and adjusts this score by a weight, $w_{i}$, proportional to the overlap’s significance to the trajectory length. Similarly, the algorithm determines $\tau_{\zeta}$ adjusted by the same weight, $w_{i}$.
Algorithm 1 Sub-task Decomposition Similarity
1:function SIMILARITY($\mathcal{S}$, $\hat{\mathcal{S}}$)
2:     $i$leftarrow 0 $\triangleright$ Number of intersecting sub-tasks
3:     for each $s_{n}$ in $\mathcal{S}$ do
4:         for each $\hat{s}_{n}$ in $\hat{\mathcal{S}}$ do
5:              if iou($s_{n},\hat{s}_{n}$) $>0$ then
6:                  $i\leftarrow i+1$
7:                  $IOU_{i}\leftarrow\textsc{iou}(s_{n},\hat{s}_{n})$
8:                  $CS_{i}\leftarrow\textsc{cosine\_similarity}(s_{n},\hat{s}_{n})$
9:                  $w_{i}\leftarrow\textsc{interval\_weight}(s_{n},\hat{s}_{n})$
10:     $\tau_{k}\leftarrow\frac{\sum{IOU_{i}}\times w_{i}}{\sum{w_{i}}}$triangleright Normalized temporal similarity
11:     $\tau_{\zeta}\leftarrow\frac{\sum{CS_{i}}\times w_{i}}{\sum{w_{i}}}$triangleright Normalized semantic similarity
12:     return $\tau_{k}$, $\tau_{\zeta}$
Ground-truth $\mathcal{S}$ $\hat{\mathcal{S}}$ using $\mathcal{D}^{kxu}$, $\textbf{C}^{1S}$ $\tau_{k}=0.87$, $\tau_{\zeta}=0.98$ $\hat{\mathcal{S}}$ using $\mathcal{D}^{kxu}$, $\textbf{C}^{0S}$ $\tau_{k}=0.74$, $\tau_{\zeta}=0.51$ Random $\tilde{\mathcal{S}}$ $\tau_{k}=0.31$, $\tau_{\zeta}=0.14$
{(0, 10, “Move to above   xxxxxxx Cube A”), (11, 23, “Move directly down   xxxxxxx to Cube A”), (24, 25, “Grasp Cube A”), (26, 39, “Vertically pick   xxxxxxx up Cube A”), (40, 48, “Align Cube A   xxxxxxx with Cube B”), (49, 54, “Move Cube A   xxxxxxx vertically down   xxxxxxx to Cube B”), (55, 58, “Release Cube A   xxxxxxx onto Cube B”), (59, 62, ‘Return Home”)} {(0, 10, “Move to above   xxxxxxx Cube A”), (11, 23, “Move directly   xxxxxxx down to Cube A”), (24, 25, “Grasp Cube A”), (26, 39, “Vertically pick   xxxxxxx up Cube A”), (40, 54, “Align Cube A   xxxxxxxwith Cube B”), (55, 58, “Release Cube A   xxxxxxx onto Cube B”), (59, 62, “Return Home’)} {(0, 10, “robot arm moving   xxxxxxx towards cube A”), (11, 24, “robot arm adjusting   xxxxxxx position above cube A”), (25, 39, “robot arm descending   xxxxxxx towards cube A”), (40, 54, “robot arm   xxxxxxx lifting cube A”), (55, 62, “robot arm   xxxxxxx moving cube A   xxxxxxx towards cube B”)} {(0, 38, “Sort objects by color”’), (39, 42, “Turn on a light”), (43, 49, “Serve a meal”), (50, 53, “Remove object from wall”), (54, 61, “Place object on a shelf”), (62, 62, “Sort objects by color”)}
TABLE I: Ground-truth $\mathcal{S}$ compared against two $\hat{\mathcal{S}}$ and one $\tilde{\mathcal{S}}$ for a single trajectory in the Stack environment: The second column shows a GPT-4V $\hat{\mathcal{S}}$ using textual data, $\mathcal{D}^{kxu}$, and an in-context example, $\textbf{C}^{1S}$. The third column shows a GPT-4V $\hat{\mathcal{S}}$ using textual data, $\mathcal{D}^{kxu}$, and excludes an in-context example, $\textbf{C}^{0S}$. The last column shows a randomly generated sub-task decomposition, $\tilde{\mathcal{S}}$. The metrics, $\tau_{k}$ and $\tau_{\zeta}$, shown at the top of each column are scores associated with $\texttt{SIMILARITY}(\mathcal{S},\hat{\mathcal{S}})$ or $\texttt{SIMILARITY}(\mathcal{S},\tilde{\mathcal{S}})$. The important note is $\hat{\mathcal{S}}$ using $\textbf{C}^{1S}$ is nearly identical to the ground-truth, while $\hat{\mathcal{S}}$ using $\textbf{C}^{0S}$ has different temporal partitions and overall different vernacular.
Example: We will explain SIMILARITY with $s_{4}$ in $\mathcal{S}$ from Eq. 2 and $\hat{s}_{3}$ in $\hat{\mathcal{S}}$ that is produced by GPT-4V using $\mathcal{D}^{kxu}$ and $\textbf{C}^{0S}$. This $\hat{\mathcal{S}}$ is fully shown in the third column of Table I. The respective sub-tasks are:
[…]
In both 5 and 7, the subroutine IOU() calculates the temporal alignment between $s_{4}$ and $\hat{s}_{3}$.
$\texttt{IOU}(s_{4},\hat{s}_{3})=\texttt{IOU}((40,48),(40,54))=8/14=0.571.$
[…]
ENCODER() represents any state-of-the-art pre-trained model capable of semantic encoding, such as USE [39]. The results above use the USE encoder and we use the USE encoder in Section V. The choice of encoder is based on its efficacy in capturing textual semantics, essential for understanding and comparing sub-task descriptions [40], [41]. This approach ensures flexibility and adaptability in applying the most suitable semantic encoding technique for the task at hand.
In 9, the subroutine INTERVAL_WEIGHT() calculates the shortest combined window length of $s_{4}$ and $\hat{s}_{3}$, normalized by maximum trajectory lengths between $K_{\mathcal{S}}$ and $K_{\hat{\mathcal{S}}}$, which both happen to be 62 in this example. The interval weight, $w_{i}$, between $s_{4}$ and $\hat{s}_{3}$ is:
[…]
$=\frac{\min(48,54)-\max(40,40)+1}{\max(62,62)}$ $=0.145.$
After all sub-task intersections are calculated, 10 and 11 calculate $\tau_{k}$ and $\tau_{\zeta}$ to be within their respective ranges.
V Experiments
Figure 3: Ablation study for two environments: Temporal, $\tau_{k}$, and semantic, $\tau_{\zeta}$, statistics for different FMs with varying input parameters including context C, and data modalities (textual data, $\mathcal{D}^{kxu}$, and visual data, $\mathcal{D}^{\mu}$). The key insight is an in-context example, $\textbf{C}^{1S}$ (green), significantly boosts both $\tau_{k}$ and $\tau_{\zeta}$, while visual data often decreases $\tau_{\zeta}$. Using our metrics across different FMs and P, developers can choose the best FM configuration for their application. In these examples GPT-4V with $\textbf{C}^{1S}$ and $\mathcal{D}^{kxu}$ performs the best and most consistent.
The goal of our experiments is to verify our approach on a large and diverse set of trajectory lengths, initial conditions, sub-task durations, and sub-task descriptions. Our experiments include 200 unique trajectories (4 environments, 50 trajectories per environment) with corresponding $\mathcal{D}$ and a ground-truth $\mathcal{S}$. All project code including experimental configurations, Algorithm 1 implementation, FSM implementations, and results generation can be found at github.com/jsalfity/task_decomposition.
All results are shown in Table II, which presents mean and standard deviation for both $\tau_{k}$ and $\tau_{\zeta}$ across 50 trajectories (50 calculations of $\texttt{SIMILARITY}(\mathcal{S},\hat{\mathcal{S}})$ for unique $\mathcal{S}$) for 2400 different combinations of FMs, including GPT-4V, Gemini Pro, Gemini Pro Vision, Video-LLaVa; input modalities, $\mathcal{D}^{kxu}$ and $\mathcal{D}^{\mu}$; and contexts, $\textbf{C}^{1S}$ and $\textbf{C}^{0S}$.
V-A Robotics’ Trajectory Data Collection
We narrow the scope of our work to robot manipulators in the Robosuite [42] simulation environment. The 4 different environments used in this work are seen in Table II including their approximate $K_{\mathcal{S}}$ and $N_{\mathcal{S}}$. The environments range from short, single objective tasks to longer, complicated multi-objective tasks. All environments are controlled using FSM-based control, seen in the supporting code, to control robot’s end-effector position based on pre-defined sub-tasks. The transition between sub-tasks is determined by spatial position conditions such as when a gripper is directly above an object, when a grasp is achieved, etc. Each sub-task is labeled directly by the FSM. As an example, the Lift environment’s sequential sub-task descriptors are: “Move to cube”, “Grasp Cube”, “Lift Cube”, and moves to the next sub-task when the previous sub-task is complete. The FSM approach enables tracking specific sub-tasks and the associated start/stop steps. In each sub-task, a position feedback controller commands the end-effector to move to the desired x-y-z position using Robosuite’s OSC_POSE controller.
The control vector is $u\in\mathbb{R}^{7}$ and includes the desired end-effector pose in $\mathbb{R}^{6}$ and a boolean value that specifies the gripper state (open or closed). The state vector, $x$, consists of the global robot state x-y-z end-effector positions in $\mathbb{R}^{3}$ and global non-robot object states, where each object is represented by x-y-z positions in $\mathbb{R}^{3}$. Examples of non-robot objects are cubes, door handles, cans, etc. The $\mathcal{D}^{\mu}$ consists of 256 $\times$ 256 RGB images, encoded into Base64 format, with the step number annotated at the top of each image, as shown in the bottom of the P in Fig. 2.
To generate a unique trajectory, environmental objects are randomly positioned. Thus the general manipulation task for each environment remain similar but each sub-task contains different start and end steps and different $K_{\mathcal{S}}$. The $N_{\mathcal{S}}$ and approximate $K_{\mathcal{S}}$ is written below each environment in Table II.
V-B Querying an FM
Four FM models are used: GPT-4V [8], Gemini Pro, Gemini Pro Vision [9], and Video-LLaVA [10]. The configuration settings for each model can be found in Section VII and the supporting code. These models offer a diversity of input modality (textual and visual) and different parametric tuning accessibility and customization. For example, Video-LLaVA is an open source model, whereas GPT-4V is only accessed by an API.
For each FM query, the P is built how it appears in Fig. 2. The only differences are $\textbf{C}^{1S}$ for each environment and $\mathcal{D}$ unique to each trajectory. The same $\textbf{C}^{1S}$ (from a hold-out $\mathcal{D}$, $\mathcal{S}$ pair) is used for each environment and consists of important transitions between sequential sub-task decompositions.
V-C Results
The experimental results, shown in Table II, are ablated across different inputs: One-Shot context, $\textbf{C}^{1S}$, Zero-Shot context, $\textbf{C}^{0S}$, textual data, $\mathcal{D}^{kxu}$, visual data, $\mathcal{D}^{\mu}$, and both textual and visual data, $\mathcal{D}$. Gemini Pro Vision and Video-LLaVa could not accept $\mathcal{D}$ due to their respective P token limit and therefore do not have such statistics.
Statistics from comparing ground-truth against a random sub-task decomposition, $\tilde{\mathcal{S}}$, is included to provide a baseline comparison. An example is shown at the right-most column of Table I and statistics are in Table II and Fig. 3. The randomized baseline for each environment is formed by $\texttt{SIMILARITY}(\mathcal{S},\tilde{\mathcal{S}})$. Each $\tilde{\mathcal{S}}$ is synthesized from a function that generates a sub-task decomposition with total length $K_{\mathcal{S}}$, but results in $N_{\tilde{\mathcal{S}}}$ random sub-tasks. Each $\tilde{s}_{n}$ is defined by a uniformly random chosen start and end step and a randomly chosen description from POSSIBLE_SUBTASKS. The function ensures there is no overlap in steps between sub-tasks. In the majority of cases, $N_{\tilde{\mathcal{S}}}\neq N_{\mathcal{S}}$, which practically means the granularity of decomposition is different. POSSIBLE_SUBTASKS contains 20 distinct $\zeta$ that are all candidates for a manipulator task to achieve. The exact function and POSSIBLE_SUBTASKS list is shown in the github repository.
Table II shows statistics over FM responses with $\hat{\mathcal{S}}$ that are valid according to Definition 3. Examples of invalid FM responses include when $k_{\text{start}}$ comes after $k_{\text{end}}$ for any given $\hat{s}_{n}$, $k_{\text{start}}$ in any $\hat{s}_{n}$ is greater than $k_{\text{start}}$ in any $\hat{s}_{n+1}$, or the FM response contains no $\hat{\mathcal{S}}$ at all. The error rates for each batch of 50 FM calls, due to 50 trajectories per environment, is reported in Table III.
It is clear from the results in Table II and the two environment examples in Fig. 3 that there is significant advantage when using $\textbf{C}^{1S}$, as all $\tau_{k}$ and $\tau_{\zeta}$ are higher compared to using $\textbf{C}^{0S}$. We speculate that this is because the in-context example directly steers the FM to attend to important transitional steps and to synthesize descriptors using similar terminology to the example.
The very low scores are consistent for $\hat{\mathcal{S}}$ that contain $N_{\hat{\mathcal{S}}}\gg N_{\mathcal{S}}$, indicating too fine of a granularity. For example, Video-LLaVa produced $N_{\hat{\mathcal{S}}}\approx K_{\mathcal{S}}$, implying a sub-task at every $k$, which is not realistic.
Open a door: Door Lift a cube: Lift Move a can: PickPlace Stack two cubes: Stack
$K_{\mathcal{S}}\approx 80$, $N_{\mathcal{S}}=5$ $K_{\mathcal{S}}\approx 40$, $N_{\mathcal{S}}=3$ $K_{\mathcal{S}}\approx 80$, $N_{\mathcal{S}}=7$ $K_{\mathcal{S}}\approx 120$, $N_{\mathcal{S}}=5$
LLM (Context, Data) $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$
GPT-4V $(\textbf{C}^{1S},\mathcal{D}^{kxu})$ $0.97\pm 0.07$ $0.87\pm 0.16$ $0.80\pm 0.15$ $0.83\pm 0.10$ $0.93\pm 0.10$ $0.60\pm 0.12$ $0.74\pm 0.12$ $0.90\pm 0.08$
GPT-4V $(\textbf{C}^{0S},\mathcal{D}^{kxu})$ $0.78\pm 0.13$ $0.30\pm 0.13$ $0.81\pm 0.16$ $0.61\pm 0.14$ $0.83\pm 0.14$ $0.45\pm 0.09$ $0.71\pm 0.11$ $0.32\pm 0.19$
GPT-4V $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ $0.70\pm 0.17$ $0.69\pm 0.18$ $0.32\pm 0.14$ $0.48\pm 0.18$ $0.25\pm 0.07$ $0.37\pm 0.06$ $0.53\pm 0.15$ $0.88\pm 0.06$
GPT-4V $(\textbf{C}^{0S},\mathcal{D}^{\mu})$ $0.44\pm 0.10$ $0.53\pm 0.07$ $0.32\pm 0.11$ $0.20\pm 0.04$ $0.25\pm 0.11$ $0.32\pm 0.07$ $0.40\pm 0.08$ $0.51\pm 0.10$
GPT-4V $(\textbf{C}^{1S},\mathcal{D})$ $0.95\pm 0.08$ $0.71\pm 0.14$ $0.78\pm 0.14$ $0.76\pm 0.05$ $0.75\pm 0.20$ $0.54\pm 0.16$ $0.70\pm 0.12$ $0.75\pm 0.09$
GPT-4V $(\textbf{C}^{0S},\mathcal{D})$ $0.77\pm 0.13$ $0.47\pm 0.17$ $0.73\pm 0.12$ $0.57\pm 0.15$ $0.57\pm 0.19$ $0.39\pm 0.09$ $0.71\pm 0.13$ $0.62\pm 0.10$
Gemini Pro $(\textbf{C}^{1S},\mathcal{D}^{kxu})$ $0.92\pm 0.06$ $0.98\pm 0.02$ $0.84\pm 0.15$ $0.52\pm 0.21$ $0.62\pm 0.13$ $0.65\pm 0.14$ $0.53\pm 0.15$ $0.88\pm 0.05$
Gemini Pro $(\textbf{C}^{0S},\mathcal{D}^{kxu})$ $0.57\pm 0.15$ $0.40\pm 0.14$ $0.64\pm 0.15$ $0.38\pm 0.09$ $0.64\pm 0.21$ $0.40\pm 0.12$ $0.48\pm 0.11$ $0.31\pm 0.10$
Gemini Pro V $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ $0.91\pm 0.07$ $0.84\pm 0.15$ $0.61\pm 0.18$ $0.74\pm 0.24$ $0.57\pm 0.19$ $0.67\pm 0.14$ $0.54\pm 0.17$ $0.89\pm 0.06$
Gemini Pro V $(\textbf{C}^{0S},\mathcal{D}^{\mu})$ $0.25\pm 0.05$ $0.21\pm 0.03$ $0.17\pm 0.04$ $0.28\pm 0.03$ $0.19\pm 0.04$ $0.38\pm 0.06$ $0.31\pm 0.04$ $0.24\pm 0.04$
Video-LLaVA $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ $0.91\pm 0.06$ $0.98\pm 0.01$ $0.58\pm 0.11$ $0.65\pm 0.02$ $0.64\pm 0.13$ $0.67\pm 0.14$ $0.32\pm 0.17$ $0.54\pm 0.01$
Video-LLaVA $(\textbf{C}^{0S},\mathcal{D}^{\mu})$ $0.08\pm 0.05$ $0.49\pm 0.04$ $0.34\pm 0.10$ $0.34\pm 0.02$ $0.11\pm 0.13$ $0.39\pm 0.07$ $0.18\pm 0.08$ $0.28\pm 0.04$
Random Baseline $0.33\pm 0.09$ $0.21\pm 0.10$ $0.43\pm 0.13$ $0.24\pm 0.04$ $0.36\pm 0.15$ $0.20\pm 0.06$ $0.32\pm 0.10$ $0.21\pm 0.06$
TABLE II: Ablation study across all environments: Comparison of temporal similarity, $\tau_{k}$, and semantic similarity, $\tau_{\zeta}$, for different robotics manipulation environments ablated across prompts, P, with different inputs modalities (textual data, $\mathcal{D}^{kxu}$ and visual data, $\mathcal{D}^{\mu}$) and context, C, (in-context example, $\textbf{C}^{1S}$, and no in-context example, $\textbf{C}^{0S}$). Statistics (mean and standard deviation) are reported over approximately 50 predictions for each environment, FM, input modality, and context combinations. The key insight is the significant increase in temporal and semantic similarities for $\hat{\mathcal{S}}$ with P that include $\mathcal{D}^{kxu}$ and $\textbf{C}^{1S}$.
VI Conclusion
We present a method to send multi-modal input data including step-by-step state, action and corresponding images to FM including LLMs and VLMs. We steer an FM for use as an analytics engine to digest raw trajectory data in the prompt and query the FM to decompose the data into distinct sub-tasks. As of this writing, the cost of offloading and requesting data is \$0.03 to \$0.15 depending on the trajectory length. We see this analysis and capability of using FMs for robotic’s data labeling as cost-effective and data efficient for the broader research problem of robot learning at scale.
The results are shown with a ground-truth $\mathcal{S}$ crafted by a customized FSM. There may be other $\mathcal{S}$ with different granularity of sub-tasks, resulting in a different number of sub-tasks, $N_{\mathcal{S}}$, and alternative descriptors. But our results are still valid because the proposed annotation pipeline and evaluation method will be set up by developers who want FM predictions, $\hat{\mathcal{S}}$, and ensuing annotations similar to $\mathcal{S}$ of their own.
The IOU calculation is a strict temporal calculation, resulting in strict weightings. This could create challenges for future developers, but an alternative “soft” IOU could be utilized where an intersection may be greater than zero for temporal ranges that are near each other, but do not necessarily intersect. In our algorithm, shorter sub-tasks are assigned smaller weights towards $\tau_{k}$ and $\tau_{\zeta}$. In practice, a short sub-task may be more important to a developer than longer sub-tasks. Users can manually increase the relative weighting of important sub-tasks to give more contribution towards the final scores. While we use the USE encoder to calculate $\tau_{\zeta}$, our framework is designed to be modular, allowing future use of any pre-trained encoder.
The presented robotics manipulation environments have structured and repeatable $\mathcal{S}$, which then allow us to present quantitative results. Due to limits on prompt tokens, longer trajectories - given the resources - can also be considered using our approach.